If a+bc=d, then which expression describes c?
A: d - a + b
B: a + b + c
C: (d-a) divided by b
D: (a+d) divided by b
: If a+bc=d, then which expression describes c?
A: d - a + b
B: a + b + c
C: (d-a) divided by b
D: (a+d) divided by b

Answer by checkley77(3848)

a+bc=d
bc=d-a
c=(d-a)/b answer=C).